OpenStudy (anonymous):

y/-2 + 5= -4 options... 2 −2 18 −18

OpenStudy (anonymous):

18

OpenStudy (anonymous):

plug em in...18/-2 = 5 = -4...18 IS CORRECT

OpenStudy (luigi0210):

multiply by -2 to get rid of the fraction

OpenStudy (andriod09):

Hey Kileyrides! Welcome to openstudy!! Let me clear them up. A)Yes, the correct answer is 18, and they should help you solve and, which leads me to b) How to solve this: the best thing to do first would be to take the fraction and multiply both sides by -2, like so: \[\frac{y}{-2} + 5= -4\] and multiply them by -2, doing so would take away the fraction: \[(-2)\frac{y}{-2}+5=-4(-2)\] doing that, would cancel out the fraction, leaving you with \[y+5=8\] then, what you would do next is simply subtract 5 from both sides, which would leave you with: \[y=18\]